Explain plank’s quantum theory?
 bohr’s theory of hydrogen atom?

Grade:11

2 Answers

Umakant biswal
7 years ago
@ sanjay 
ask single ques in one thread , so, that it will be easy to explain us . 
bohr suggested a proper model for hydrogen like atoms , 
here are the some of the imp colclusions from his model 
1- acc to him , electrons are orbiting around the nucleaus , in a circular orbit  
2- but only those orbits are allowed whoose , angular momentum is the integral multiple of h/ 2[pie 
3- bohr called these as the quantised orbit and acc to him electron will not loose energy if it will move in that orbit . 
4- energy will be required if electron will make transition from one energy level to anathor energy level . 
5- high energy states are not stable and lectron will be back to its ground state after very short period of time . 
and if theleectron will gain energy then its called as absorption spectra , and if it will loose some amount of energy its called as emission spectra . 
HOPE IT CLEARS 
Vikas TU
7 years ago
According to plancks qunatum theory the photons striking the photosensitive elements/matreial would emit electrons only and only if the striking energy is greater than its threshold energy and thus photoelctric effect takes place according to the equation.
that is:
 
E = W + K.E
hv = hv0 + eV
 
ANd Bohr’s theory explain about the atoms taking hydrogen as the reference that electrons are revolving in an robit around the nucleus and doesn’t falls into it.
and gave some postulates that is folowed for H-like atoms.
For exapmle the angular momentum would be equal to =>
mvr =nh/2pi
and radius of ato m = 0.529*-n^2/z
  • and Energy = -13.6z^2/n^2